OpenStudy (anonymous):

After 43.0 minutes, 34.0% of compound decomposed. What is the half-life of this reaction assuming first order kinetics?

OpenStudy (anonymous):

let for first order kinetics; \[A_0 = initial _ . amount\] \[A_t = remaining _. amount _. at _. time_. t\] k = reaction rate coefficient \[\ln[A]_t = -kt + \ln[A]_0\] \[\ln66A_0 = -43k + lnA_0\] now we can find k. for half-life:\[t/2 = \ln2/k\]

OpenStudy (anonymous):

whats the value of k?

OpenStudy (anonymous):

anyone..?